OpenStudy (anonymous):

(Medal given to first complete answer) Write and factor a problem that is a quadratic expression, then check your solutions by multiplying after you factor.

OpenStudy (anonymous):

y=x^2+x-6 y=(x+3)(x-2) y=x^2-2x+3x-6 y=x^2+x-6

OpenStudy (anonymous):

Mind explaining it some so I can get a better understanding?